OpenStudy (anonymous):

A student has an average score of 85% after completing 4 tests. What is the lowest possible score in any one of the tests?

OpenStudy (anonymous):

Is this one really difficult for people? or does no one want to help?

OpenStudy (australopithecus):

My reasoning is that for him to get his lowest possible score all the other tests would have to be perfect. So thus you have the equation: (1+1+1+x)/4 = 0.85

OpenStudy (australopithecus):

(3+x)/4 = 0.85 3 + x = 4(0.85) x = 4(0.85) - 3 x = 0.40 thus He the lowest possible score he could have gotten on one of his tests is a 40% Just because no one is helping you doesn't mean the question is difficult just be patient or ask in chat for help

OpenStudy (australopithecus):

1 is a perfect score btw or 100%